OpenStudy (anonymous):

Graph the line. -2x+y=-6

OpenStudy (solomonzelman):

-2x+y=-6 y=2x-6 plug in any 2 random numbers for x x=2 y=2(2)-6 y=-2 so you 1st point is (2,-2) x=1 y=2(1)-6 y=-4 so you 2nd point is (1,-4) or you can use the y-intercept which is -6, or as coordinate (0,-6) connect these points (any 2) with a line.